# Microwave-to-Optical Quantum Transduction with Antiferromagnets

### v1

@@ VI Estimation of the Transduction Efficiency
It should be noted that the microwave-magnon interaction strength $g_{\mu}$ [Eq. (10)] depends explicitly on the frequency of the microwave cavity, i.e., the antiferromagnetic resonance frequency as $\propto\sqrt{\omega_{\mathrm{e}}}$, while the light-magnon interaction strength $\zeta_{\mu}$ [Eq. (12)] does not. We take as an example $g_{0}/2\pi=A\sqrt{\omega_{\mathrm{e}}}$ with $A=25\,\mathrm{mHz}/\sqrt{1\,\mathrm{GHz}}$ [35]. We also assume that $G_{\mu}/2\pi=0.1\kappa_{\mu}/\sqrt{10^{9}V_{\mathrm{AFM}}/[\mathrm{mm^{3}}]}\,\mathrm{MHz}$ where $V_{\mathrm{AFM}}$ is given in units of mm3 [36].
@@ VI.1 The case with an optical cavity
where $\mu=\alpha$ or $\beta$, $\eta_{\mathrm{o}}=\kappa_{\mathrm{o,e}}/\kappa_{\mathrm{o}}$, $\eta_{\mathrm{e}}=\kappa_{\mathrm{e,e}}/\kappa_{\mathrm{e}}$, and $C_{\mathrm{em},\mu}=4g_{\mu}^{2}/(\kappa_{\mathrm{e}}\gamma_{\mu})$ and $C_{\mathrm{om},\mu}=4\zeta_{\mu}^{2}/(\kappa_{\mathrm{o}}\gamma_{\mu})$ are the cooperativity between microwave photon and magnon and the cooperativity between optical photon and magnon, respectively. Eq. (39) coincides with a generic expression for the transduction efficiency mediated by one intermediate bosonic mode [6]. To estimate the magnitude of the transduction efficiency from Eq. (39), suppose that the lower-frequency magnon mode with $\omega_{\beta}/2\pi=20\,\mathrm{GHz}$ is excited in an easy-axis antiferromagnet, as experimentally realized [35]. We use the following possible values of parameters: $g_{\beta}/2\pi\approx 3.3\,\mathrm{MHz}$, $\zeta_{\beta}/2\pi\approx 40\,\mathrm{kHz}$ (with $\kappa_{\beta}\approx 0.4$ and $\bar{n}_{\mathrm{cav}}\approx 1\times 10^{6}$), $\gamma_{\beta}/2\pi\approx 100\,\mathrm{MHz}$, $\kappa_{\mathrm{o,i}}/2\pi=\kappa_{\mathrm{o,e}}/2\pi\approx 100\,\mathrm{MHz}$, and $\kappa_{\mathrm{e,i}}/2\pi=\kappa_{\mathrm{e,e}}/2\pi\approx 100\,\mathrm{MHz}$ [29, 36, 35]. Substituting these possible values into Eq. (39), the magnitude of the transduction efficiency is estimated to be
@@ VI.3 Thickness dependence
As shown in Fig. 4, we find a distinct behavior of the transduction efficiency $\eta$ as a function of the thickness $d_{\mathrm{AFM}}$ depending on the presence or absence of an optical cavity. First, we show the result for the case with an optical cavity in Fig. 4(a). In this case, we see that there exists an optimal value of the thickness $d_{\mathrm{AFM}}\sim\mathcal{O}(10^{-3}\,\mathrm{mm})$ at which the transduction efficiency takes a maximum value. This result can be understood as follows. When the value of $d_{\mathrm{AFM}}$ is small, the effect of the light-magnon interaction $\zeta_{\beta}\propto 1/\sqrt{d_{\mathrm{AFM}}}$, which is an decreasing function of $d_{\mathrm{AFM}}$, is dominant, and thus the transduction efficiency scales as $\eta\sim C_{\mathrm{om},\beta}^{-1}\sim d_{\mathrm{AFM}}$. On the other hand, when the value of $d_{\mathrm{AFM}}$ is large, the effect of the microwave-magnon interaction $\xi_{\beta}\propto\sqrt{d_{\mathrm{AFM}}}$, which is an increasing function of $d_{\mathrm{AFM}}$, is dominant, and thus the transduction efficiency scales as $\eta\sim C_{\mathrm{em},\beta}^{-1}\sim d_{\mathrm{AFM}}^{-1}$.
Next, we show the result for the case without an optical cavity in Fig. 4(b). In this case, we see that the transduction efficiency is a monotonically increasing function of the thickness $d_{\mathrm{AFM}}$. This result can be understood from that both the light-magnon interaction $\xi_{\beta}$ and the microwave-magnon interaction $g_{\beta}$ are increasing functions of $d_{\mathrm{AFM}}$, and thus the transduction is also an increasing function of $d_{\mathrm{AFM}}$.
@@ VII Discussion
The magnitudes of the transduction efficiency we have estimated so far are not sufficient for practical use as a quantum transducer, since it has been suggested theoretically that $\eta=0.5$ is the lower limit for realizing the quantum state transfer between quantum devices [4, 5, 6]. Let us consider possible ways of improving the transduction efficiency. As we have seen in Secs. VI.1 and VI.2, in both cases with and without an optical cavity the transduction efficiency gets improved as the magnon decay rate becomes smaller. Thus, utilizing antiferromagnets with a low magnon decay rate is one possible direction. Another direction may be to utilize the heterostructures consisting of an antiferromagnet and a nonmagnetic insulator, in a similar way to the topological insulator heterostructures [30]. Here, we take a look at the case with an optical cavity, because there is a sample size limitation in the case without an optical cavity where there is no limitation in the case with an optical cavity (see Sec. IV.2). For a heterostructure of $N_{\mathrm{L}}$ antiferromagnet layers and $N_{\mathrm{L}}$ nonmagnetic insulator layers, the microwave-magnon and light-magnon interaction strengths are modified as $g_{\mu}\to\tilde{g}_{\mu}=g_{\mu}\sqrt{N_{\mathrm{L}}}$ and $\zeta_{\mu}\to\tilde{\zeta}_{\mu}=\zeta_{\mu}\sqrt{N_{\mathrm{L}}}$, respectively. This is because the magnon operators are modified as $\hat{m}_{\mu}\to\hat{\mathfrak{m}}_{\mu}=\frac{1}{\sqrt{N_{\mathrm{L}}}}\sum_{i=1}^{N_{\mathrm{L}}}\hat{m}_{i,\mu}$, where $\hat{m}_{i,\mu}$ is the magnon operator of the $i$-th antiferromagnet layer and $\hat{\mathfrak{m}}_{\mu}$ is the collective magnon operator satisfying the commutation relation $[\hat{\mathfrak{m}}_{\mu},\hat{\mathfrak{m}}_{\mu}^{\dagger}]=1$ [30]. Then, we find that the transduction efficiency $\eta$ [Eq. (39)] in the triple resonance condition scales as $\eta\sim N_{\mathrm{L}}^{2}$, since with the parameters we have used in this study the cooperativities are small so that $\eta\sim C_{\mathrm{om},\mu}C_{\mathrm{em},\mu}$ in Eq. (39). If a heterostructure with $N_{\mathrm{L}}=5000$ and $d_{\mathrm{AFM}}=1\,\mathrm{\mu m}$ (the thickness of each layer) is possible, then the transduction efficiency is improved from $\eta\sim 10^{-9}$ [Eq. (40)] to
@@
Finally, we mention a recent study on a microwave-to-optical quantum transduction using antiferromagnetic topological insulators [46]. The quantum transduction studied in Ref. [46] is based on a strong nonlinear optical interaction via spin-orbit coupling, which gives rise to the effective Hamiltonian of the form $H_{\mathrm{int}}=\hbar\mathcal{G}(\hat{a}^{\dagger}\hat{b}+\hat{b}^{\dagger}\hat{a})$, where $\hat{a}$ and $\hat{b}$ are the annihilation operator of microwave and optical photons, respectively. This mechanism is similar to that of the quantum transduction via the electro-optic effect, which is thus different from ours, i.e., the quantum transduction via the intermediate bosonic mode.
@@ VIII Summary
To summarize, we have formulated a theory for the microwave-to-optical quantum transduction mediated by antiferromagnetic magnons in antiferromagnets. We have derived analytical expressions for the transduction efficiency in the cases with and without an optical cavity, where a microwave cavity is used in both cases. The derived expressions for the transduction efficiency take complicated forms when the two antiferromagnetic magnon modes coexist, while they reduce to the same expressions as in the case utilizing ferromagnets when one of the two antiferromagnetic magnon modes are excited. We have found that the quantum transduction can occur even in the absence of an external static magnetic field, which is in contrast to the case of the quantum transduction using ferromagnets where an external static magnetic field is essential to cause the resonance state. We have also found that, in the case with an optical cavity the transduction efficiency takes a peak structure with respect to the sample thickness, indicating that there exists an optimal thickness, whereas in the case without an optical cavity the transduction efficiency is a monotonically increasing function of the sample thickness. It is expected that by utilizing a heterostructure consisting of an antiferromagnet and a nonmagnetic insulator the transduction efficiency can be improved to the magnitude that is comparable to the up-to-date values achieved so far in optomechanical and electro-optic systems. Owing to the wide tunability of the antiferromagnetic resonance frequency, it is expected that a variety of quantum devices that operate at the wide frequency ranges from $\mathcal{O}(1\,\mathrm{GHz})$ to $\mathcal{O}(1\,\mathrm{THz})$ can be interconnected via the quantum transducer utilizing antiferromagnets. Our study opens up a way for possible applications of antiferromagnetic materials in future quantum interconnects.

### v2

@@ VI Estimation of the Transduction Efficiency
It should be noted that the microwave-magnon interaction strength $g_{\mu}$ [Eq. (10)] depends explicitly on the frequency of the microwave cavity, i.e., the antiferromagnetic resonance frequency as $\propto\sqrt{\omega_{\mathrm{e}}}$, while the light-magnon interaction strength $\zeta_{\mu}$ [Eq. (12)] does not. We take as an example $g_{0}/2\pi=A\sqrt{\omega_{\mathrm{e}}}$ with $A=25\,\mathrm{mHz}/\sqrt{1\,\mathrm{GHz}}$ [35]. Unless otherwise noted, we also assume that $G_{\mu}/2\pi=0.1\kappa_{\mu}/\sqrt{10^{9}V_{\mathrm{AFM}}/[\mathrm{mm^{3}}]}\,\mathrm{MHz}$ [36] with $V_{\mathrm{AFM}}$ being given in units of mm3, which is obtained from the values for YIG [48]: $\theta_{\mathrm{F}}=\theta_{\mathrm{F,YIG}}\approx 20^{\circ}\mathrm{/mm}$, the spin density $n_{\mathrm{YIG}}=2.1\times 10^{19}\,\mathrm{mm^{-3}}$, and $\varepsilon_{\mathrm{r}}\approx 5$.
@@ VI.1 The case with an optical cavity
where $\mu=\alpha$ or $\beta$, $\eta_{\mathrm{o}}=\kappa_{\mathrm{o,e}}/\kappa_{\mathrm{o}}$, $\eta_{\mathrm{e}}=\kappa_{\mathrm{e,e}}/\kappa_{\mathrm{e}}$, and $C_{\mathrm{em},\mu}=4g_{\mu}^{2}/(\kappa_{\mathrm{e}}\gamma_{\mu})$ and $C_{\mathrm{om},\mu}=4\zeta_{\mu}^{2}/(\kappa_{\mathrm{o}}\gamma_{\mu})$ are the cooperativity between microwave photons and antiferromagnetic magnons and the cooperativity between optical photons and antiferromagnetic magnons, respectively. Eq. (39) coincides with a generic expression for the transduction efficiency mediated by one intermediate bosonic mode [6]. To estimate the magnitude of the transduction efficiency from Eq. (39), suppose that the lower-frequency magnon mode with $\omega_{\beta}/2\pi=20\,\mathrm{GHz}$ is excited in an easy-axis antiferromagnet [^3] Note that in the case of easy-axis antiferromagnets a relatively high magnetic field of $\sim 5\,\mathrm{T}$ would be required for realizing a lower-frequency mode of $\sim 20\,\mathrm{GHz}$. On the other hand, in a canted easy-plane antiferromagnet with the Dzyaloshinskii-Moriya interaction, a low magnetic field of $\sim 0.5\,\mathrm{T}$ can be sufficient to realize such a lower-frequency mode [35]., as experimentally realized [35]. We use the following possible values of parameters: $g_{\beta}/2\pi\approx 3.3\,\mathrm{MHz}$, $\zeta_{\beta}/2\pi\approx 40\,\mathrm{kHz}$ (with $\kappa_{\beta}\approx 0.4$ and $\bar{n}_{\mathrm{cav}}\approx 1\times 10^{6}$), $\gamma_{\beta}/2\pi\approx 100\,\mathrm{MHz}$, $\kappa_{\mathrm{o,i}}/2\pi=\kappa_{\mathrm{o,e}}/2\pi\approx 100\,\mathrm{MHz}$, and $\kappa_{\mathrm{e,i}}/2\pi=\kappa_{\mathrm{e,e}}/2\pi\approx 100\,\mathrm{MHz}$ [29, 36, 35]. Substituting these possible values into Eq. (39), the magnitude of the transduction efficiency is estimated to be
@@
Figure 4: Transduction efficiency $\eta$ as a function of the Faraday rotation angle per sublattice (per unit length) $\theta_{\mathrm{F}}$ in the case with an optical cavity. Here, $\theta_{\mathrm{F,YIG}}$ is the Faraday rotation angle per unit length of the ferromagnet YIG. Equation (40) corresponds to the value with $\theta_{\mathrm{F}}/\theta_{\mathrm{F,YIG}}=1$ in this figure.
In the process of estimating the magnitude of the transduction efficiency [Eq. (40)], we have assumed that $\theta_{\mathrm{F}}/\theta_{\mathrm{F,YIG}}=1$ due to the lack of the experimental data on the value of $\theta_{\mathrm{F}}$ (the Faraday rotation angle per sublattice) in antiferromagnets. However, this assumption might be inappropriate, since in general the magnitude of the one-magnon Raman scattering is rather small compared to that of the two-magnon Raman scattering in antiferromagnets [41, 42]. Therefore, we plot in Fig. 4 the transduction efficiency $\eta$ as a function of $\theta_{\mathrm{F}}$ (or equivalently, as a function of $G_{\beta}$). We find that the transduction efficiency monotonically increases as the value of $\theta_{\mathrm{F}}$ increases. When $\theta_{\mathrm{F}}/\theta_{\mathrm{F,YIG}}=0.01$, we have $\eta\sim 10^{-13}$. Because $C_{\mathrm{om},\mu}\ll 1$ in our setup, it follows from Eq. (39) that the tranduction efficiency scales as $\eta\sim\zeta_{\beta}^{2}\propto\theta_{\mathrm{F}}^{2}$.
@@ VI.3 Thickness dependence
As shown in Fig. 5, we find a distinct behavior of the transduction efficiency $\eta$ as a function of the thickness $d_{\mathrm{AFM}}$ depending on the presence or absence of an optical cavity. First, we show in Fig. 5(a) the result for the case with an optical cavity, which is obtained from Eq. (39). In this case, we see that there exists an optimal value of the thickness $d_{\mathrm{AFM}}\sim\mathcal{O}(10^{-3}\,\mathrm{mm})$ at which the transduction efficiency takes a maximum value. This result can be understood as follows. When the value of $d_{\mathrm{AFM}}$ is small, the effect of the light-magnon interaction $\zeta_{\beta}\propto 1/\sqrt{d_{\mathrm{AFM}}}$, which is an decreasing function of $d_{\mathrm{AFM}}$, is dominant, and thus the transduction efficiency scales as $\eta\sim C_{\mathrm{om},\beta}^{-1}\sim d_{\mathrm{AFM}}$. On the other hand, when the value of $d_{\mathrm{AFM}}$ is large, the effect of the microwave-magnon interaction $\xi_{\beta}\propto\sqrt{d_{\mathrm{AFM}}}$, which is an increasing function of $d_{\mathrm{AFM}}$, is dominant, and thus the transduction efficiency scales as $\eta\sim C_{\mathrm{em},\beta}^{-1}\sim d_{\mathrm{AFM}}^{-1}$.
Next, we show in Fig. 5(b) the result for the case without an optical cavity, which is obtained from Eq. (42). In this case, we see that the transduction efficiency is a monotonically increasing function of the thickness $d_{\mathrm{AFM}}$. This result can be understood from that both the light-magnon interaction $\xi_{\beta}$ and the microwave-magnon interaction $g_{\beta}$ are increasing functions of $d_{\mathrm{AFM}}$, and thus the transduction is also an increasing function of $d_{\mathrm{AFM}}$.
@@ VII Discussion
[…]
@@
Figure 7: Schematic illustration of a heterostructure consisting of $N_{\mathrm{L}}$ antiferromagnetic insulator layers and $N_{\mathrm{L}}$ nonmagnetic insulator layers.
[…]
@@ VIII Summary
To summarize, we have formulated a theory for the microwave-to-optical quantum transduction mediated by antiferromagnetic magnons in antiferromagnets. We have derived analytical expressions for the transduction efficiency in the cases with and without an optical cavity, where a microwave cavity is used in both cases. The derived expressions for the transduction efficiency [Eqs. (30) and (37)] take complicated forms when the two antiferromagnetic magnon modes coexist, while they reduce to the same expressions as in the case utilizing ferromagnets when one of the two antiferromagnetic magnon modes are excited. Here, we stress that the derived expressions for the transduction efficiency [Eqs. (30) and (37)] cannot be written simply as the sum of the transduction efficiency of each antiferromagnetic mode, which implies that the extension of the ferromagnetic case to the antiferromagnetic case is nontrivial. We have found that the quantum transduction can occur even in the absence of an external static magnetic field, which is in contrast to the case of the quantum transduction using ferromagnets where an external static magnetic field is essential to cause the resonance state. We have also found that, in the case with an optical cavity the transduction efficiency takes a peak structure with respect to the sample thickness, indicating that there exists an optimal thickness, whereas in the case without an optical cavity the transduction efficiency is a monotonically increasing function of the sample thickness. It is expected that by utilizing a heterostructure consisting of an antiferromagnet and a nonmagnetic insulator the transduction efficiency can be improved to the magnitude that is comparable to the up-to-date values achieved so far in optomechanical and electro-optic systems. Owing to the wide tunability of the antiferromagnetic resonance frequency, it is expected that a variety of quantum devices that operate at the wide frequency ranges from $\mathcal{O}(1\,\mathrm{GHz})$ to $\mathcal{O}(1\,\mathrm{THz})$ can be interconnected via the quantum transducer utilizing antiferromagnets. Our study opens up a way for possible applications of antiferromagnetic materials in future quantum interconnects.
@@ Acknowledgements.
Appendix A Derivation of the light-magnon interaction in antiferromagnets
In this Appendix, we outline the derivation of the light-magnon interaction in antiferromagnets [Eqs. (11) and (12)], following Refs. [36, 51, 52]. The Hamiltonian describing the interaction between light and magnon comes from the electromagnetic energy $U=\int d^{3}r\,\bm{E}\cdot\bm{D}/2$, where $\bm{D}=\hat{\varepsilon}\bm{E}$ with $\hat{\varepsilon}_{ij}$ the permittivity tensor and $\bm{E}$ an external electric field. Since the permittivity tensor can be spatial dependent, we have in general [53]
$H_{\mathrm{int}}$ $=\frac{1}{4}\int d^{3}r\,\sum_{\mu,\nu}E^{*}_{\mu}(\bm{r})\varepsilon_{\mu\nu}(\bm{r})E_{\mu}(\bm{r})$
$=\frac{V}{4N}\sum_{i}\sum_{\mu,\nu}E^{*}_{\mu}(\bm{r}_{i})\varepsilon_{\mu\nu}(\bm{r}_{i})E_{\mu}(\bm{r}_{i}),$ (46)
where $\mu,\nu=x,y,z$ denote a spatial direction, $V$ is the system volume, $N$ is the total number of lattice sites, and $i$ runs over the all lattice cites. In magnetic systems, the permittivity tensor can be expanded in terms of spins on lattice cites. Up to the terms quadratic in spin operators, we get [54, 51, 52]
$\varepsilon_{\mu\nu}(\bm{r}_{i})=\sum_{\rho}K_{\mu\nu\rho}S^{\rho}_{i}+\sum_{\alpha,\beta}G_{\mu\nu\alpha\beta}S^{\alpha}_{i}S^{\beta}_{i}+\sum_{\alpha,\beta}B_{\mu\nu\alpha\beta}S^{\alpha}_{i}S^{\beta}_{i+\delta}.$ (47)
For antiferromagnets with a rutile crystal structure such as FeF2 and MnF2 which we are interested in, the third term in the right-hand side is very small such that $B_{\mu\nu\alpha\beta}/G_{\mu\nu\alpha\beta}\sim J/\lambda\ll 1$ [54], where $J$ is the exchange coupling strength and $\lambda$ is the spin-orbit coupling strength.
In the present study, we are focusing on the one-magnon scattering process in antiferromagnets. For this purpose, the second term ($\propto S^{z}S^{\pm}$) in the right-hand side of Eq. (47) can be absorbed into the first term ($\propto S^{\pm}$). Thus, in the following, we shall take into account only the first term in Eq. (47) when deriving the light-magnon interaction.
A.1 The case of ferromagnets
For comparison, let us firstly consider the light-magnon interaction in ferromagnets. In the case of ferromagnets, which can be characterized by their macroscopic magnetization $\bm{M}$, the permittivity tensor is written as a function of the magnetization. To linear order in the magnetization, we have [see Eq. (47)] [53]
$\varepsilon_{\mu\nu}(\bm{M})=\varepsilon_{0}\left(\varepsilon_{\mathrm{r}}\delta_{\mu\nu}-iK_{0}\sum_{\rho}\epsilon_{\mu\nu\rho}M_{\rho}\right),$ (48)
where $\varepsilon_{0}$ ($\varepsilon_{\mathrm{r}}$) is the vacuum (relative) permittivity, $\epsilon_{\mu\nu\rho}$ is the Levi-Civita symbol, and $K_{0}$ is a material-dependent constant. Then, from Eqs. (46) and (48), we obtain
$H_{\mathrm{int}}=-i\frac{\varepsilon_{0}K_{0}}{4}\int d^{3}r\,\bm{M}(\bm{r})\cdot\left[\bm{E}^{*}(\bm{r})\times\bm{E}(\bm{r})\right].$ (49)
The electric field in an optical cavity can be quantized as $\bm{E}(\bm{r},t)=\sum_{\alpha}\bm{E}_{\alpha}(\bm{r})\hat{b}_{\alpha}(t)$, where $\bm{E}_{\alpha}(\bm{r})$ is the $\alpha$-th eigenmode of the electric field and $\hat{b}_{\alpha}(t)$ is the annihilation operator of the photon in the $\alpha$-th eigenmode.
For concreteness, suppose that the magnetization in the ground state is along the $z$ axis and the magnetic moments are precessing around the $z$ axis in the ferromagnetic resonance state. In this case, the electric field propagating in the direction perpendicular to the $z$ axis interact with the fluctuating component of the magnetic moments, i.e., the ferromagnetic magnon. We also have the relation $M_{i}/M_{\mathrm{s}}=\hat{S}_{i}/(SN)$ with $M_{\mathrm{s}}$ being the saturation magnetization and $SN$ being the total spin number. Finally, we obtain the light-magnon interaction in ferromagnets [49]
$H_{\mathrm{int}}=\hbar G_{\mathrm{FM}}\left(\hat{m}+\hat{m}^{\dagger}\right)\left(\hat{b}_{R}^{\dagger}\hat{b}_{R}-\hat{b}_{L}^{\dagger}\hat{b}_{L}\right),$ (50)
where $\hat{m}$ is the annihilation operator for the ferromagnetic magnon mode. The coupling strength is given by
$G_{\mathrm{FM}}=\frac{c\theta_{\mathrm{F}}}{4\sqrt{\varepsilon_{\mathrm{r}}}}\frac{1}{\sqrt{2SN}},$ (51)
with the Faraday rotation angle per unit length $\theta_{\mathrm{F}}=\omega_{\mathrm{c}}K_{0}M_{\mathrm{s}}/(2c\sqrt{\varepsilon_{\mathrm{r}}})$ ($\omega_{\mathrm{c}}$ is the cavity mode frequency).
A.2 The case of antiferromagnets
Now, let us consider the case of antiferromagnets. For concreteness, we focus on the antiferromagnets with a rutile crystal structure such as FeF2, MnF2, and CoF2 [51, 47]. In these antiferromagnets, the antiferromagnetic spins are aligned along the $z$ (crystal $c$) axis in the ground states. The magnetic ions form a body-centered tetragonal structure, with each magnetic ion surrounded by six non-magnetic ions that form a distorted octahedron. The two sublattices $A$ and $B$ (the one occupying the body-centered sites and the one occupying the corner sites) are distinguished by a 90 rotation about the $z$ axis. We choose $x$, $y$, and $z$ axes so that the magnetic ions have a two-fold rotational symmetry with respect to each axis. Then, from a point-group symmetry analysis, it turns out that for sublattice $A$ [51]
$K_{yzx}^{A}=-K_{zyx}^{A}\equiv K_{1},$
$K_{zxy}^{A}=-K_{xzy}^{A}\equiv K_{2},$
$K_{xyz}^{A}=-K_{yxz}^{A}\equiv K_{3}.$ (52)
Since sublattice $B$ is related to sublattice $A$ by a 90 rotation about the $z$ axis, we have the relation $K_{\mu\nu\rho}^{B}=\sum_{\mu^{\prime},\nu^{\prime},\rho^{\prime}}R_{\mu\mu^{\prime}}R_{\nu\nu^{\prime}}R_{\rho\rho^{\prime}}K_{\mu^{\prime}\nu^{\prime}\rho^{\prime}}^{A}$, where $R_{\alpha\beta}$ is a matrix element of the 90 rotation matrix about the $z$ axis. Therefore, for sublattice $B$,
$K_{yzx}^{B}=-K_{zyx}^{B}=-K_{xzy}^{A}=K_{2},$
$K_{zxy}^{B}=-K_{xzy}^{B}=-K_{zyx}^{A}=K_{1},$
$K_{xyz}^{B}=-K_{yxz}^{B}=-K_{yxz}^{A}=K_{3}.$ (53)
Note that $K_{1}$, $K_{2}$, and $K_{3}$ are purely imaginary.
As we have discussed briefly, we are focusing on the one-magnon scattering process. This means that the spin operators $S^{x}_{i}$ and $S^{y}_{i}$ are relevant, while $S^{z}_{i}$ is not because $S^{z}_{i}$ is quadratic in the magnon operators $\hat{a}$ and $\hat{b}$. Defining $K_{\pm}=i(K_{1}\pm K_{2})/4$, Eq. (46) reduces to [36, 51, 52]
$H_{\mathrm{int}}=$ $\ \frac{K_{+}V}{4N}\sum_{i\in A,B}\left(P_{i}^{+}S_{i}^{-}-P_{i}^{-}S_{i}^{+}\right)$
$+\frac{K_{-}V}{4N}\left[\sum_{i\in A}\left(P_{i}^{+}S_{i}^{+}-P_{i}^{-}S_{i}^{-}\right)-\sum_{j\in B}\left(P_{j}^{+}S_{j}^{+}-P_{j}^{-}S_{j}^{-}\right)\right],$ (54)
where $P_{i}^{\pm}=E^{*}_{z}(\bm{r}_{i})E_{\pm}(\bm{r}_{i})-E^{*}_{\mp}(\bm{r}_{i})E_{z}(\bm{r}_{i})$ with $E_{\pm}=E_{x}\pm iE_{y}$. Finally, by substituting an expression for the quantized electric field in an optical cavity into Eq. (54), we arrive at the expression for the light-magnon interaction, which corresponds to the one-magnon scattering, in the main text [Eqs. (11) and (16)]:
$H_{\zeta}=-\hbar\left(\hat{b}_{R}^{\dagger}\hat{b}_{R}-\hat{b}_{L}^{\dagger}\hat{b}_{L}\right)\left[G_{\alpha}\left(\hat{m}_{\alpha}^{\dagger}+\hat{m}_{\alpha}\right)+G_{\beta}\left(\hat{m}_{\beta}^{\dagger}+\hat{m}_{\beta}\right)\right].$ (55)
Here, the coupling strength is given by
$G_{\mu}=\frac{c\theta_{\mathrm{F}}}{4\sqrt{\varepsilon_{\mathrm{r}}}}\frac{\kappa_{\mu}}{\sqrt{2SN}},$ (56)
with the Faraday rotation angle per unit length $\theta_{\mathrm{F}}=\omega_{\mathrm{c}}K_{+}S/(c\sqrt{\varepsilon_{\mathrm{r}}})$. Since $K_{+}$ (i.e., $K_{1}$ and $K_{2}$) is defined as a quantity characterizing the contribution from each sublattice, the quantity $\theta_{\mathrm{F}}$ may be called the Faraday rotation angle per sublattice, or more generally, the one-magnon scattering coefficient.